if you are 3/7 of the way and have 84 miles left to drive, how long is the entire length of the trip

Answer by lwsshak3(11628)

if you are 3/7 of the way and have 84 miles left to drive, how long is the entire length of the trip
let x=entire length of trip
3/7 of the total trip already completed, so this leaves 4/7 of the entire trip yet to be completed
(4/7)x=84
4x=7*84
x=7*21=147
entire length of trip=147 miles